ClassixTailsBeatz:

In the formula d = rt, d is isolated. You can "rearrange" a formula to isolate any variable by using inverse operations. This is called solving for a variable. Solve the equation for r: Solve the equation for t:

snowflake0531:

d=rt Do you know how to solve for variables? if this was a regular equation with only one variable?

ClassixTailsBeatz:

It's confusing me tho

snowflake0531:

d=rt Let's first try to solve for r Can you divide t from each side of d=rt to isolate teh variable r?

ClassixTailsBeatz:

I don't think so

snowflake0531:

Look at it as\[d=rt \] and then divide t from each side to isolate teh r\[d/t = rt/t\]

snowflake0531:

You can split rt/t to \[\frac{ r }{ t } * \frac{ t }{ t }\] what is t/t

ClassixTailsBeatz:

I think it cancels each other out

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 You can split rt/t to \[\frac{ r }{ t } * \frac{ t }{ t }\] what is t/t \(\color{#0cbb34}{\text{End of Quote}}\) I meant d/t

snowflake0531:

So it would simplify to d/t = r

ClassixTailsBeatz:

Oh Okay, thx

ClassixTailsBeatz:

I get it now

snowflake0531:

So you got t too then?

ClassixTailsBeatz:

Yes

snowflake0531:

kk, nice

ClassixTailsBeatz:

Thx again for all your help! Really appreciate it!

snowflake0531:

yw~